OpenStudy (anonymous):

What is the rate of a first-order reaction that has a reactant concentration of 0.1M and a rate constant of 0.1/s?

OpenStudy (frostbite):

The rate of a first order reaction: \[\large v = -\frac{ d \left[ \sf{S} \right] }{ dt }=k \times \left[ S \right]\] Calculate the rate.

OpenStudy (jfraser):

Since the reaction is first order, the rate is:\[Rate = k*[Conc]^1\]You have k and the concentration, so solve for rate